A cruise boat travels 48 miles downstream in 3 hours and returns upstream in 6 hours. Find the rate of the stream.
Answer by checkley71(8403)

48/3=16 mph going down stream.
48/6=8 mph going up stream.
16-8=8 mph difference
8/2=4 mph for the current.